OpenStudy (anonymous):

Basic multivariable question. div(f(x,y,z))=∇⋅f(x,y,z), for f(x,y,z)=f(x)i+f(y)j+f(z)k is ∇⋅f(x,y,z)=((∂/∂x)f(x)i+(∂/∂y)f(y)j+(∂/∂z)f(z)k)⋅(f(x)i+f(y)j+f(z)k) ∇⋅f(x,y,z)=∇f(x,y,z)cos(φ)=((((∂/∂x)f(x))^2+((∂/∂y)f(y))^2+((∂/∂z)f(z)k)^2)^(1/2))((f(x)^2+f(y)^2+f(z)^2)^(1/2))cos(φ) But ∇⋅f(x,y,z)=(∂/∂x)f(x)i+(∂/∂y)f(y)j+(∂/∂z)f(z)k according to textbook... I'm confused...

OpenStudy (turingtest):

∇⋅f(x,y,z)=(∂/∂x)f(x)i+(∂/∂y)f(y)j+(∂/∂z)f(z)k is how I know it where did you see the other one?

OpenStudy (anonymous):

Isn't it according to the rules of dot product? a⋅b=abcos(φ)

OpenStudy (turingtest):

actually I know it as\[\nabla \cdot f=\frac{\partial f}{\partial x}+\frac{\partial f}{\partial y}+\frac{\partial f}{\partial z}\]remember that del is not a function, it is an /operator/ like d/dx

OpenStudy (anonymous):

Oh, it's an operator. Gah, why do I keep forgetting stuff like this. This is why I need to review. X( But isn't the gradient the vector sum, not just the sum of magnitudes?

OpenStudy (anonymous):

Oh, right, the derivative of the function f will include the unit vectors.

OpenStudy (turingtest):

\[\nabla\neq\frac{\partial f}{\partial x}+\frac{\partial f}{\partial y}+\frac{\partial f}{\partial z}\]but\[\nabla \cdot f=\frac{\partial f}{\partial x}+\frac{\partial f}{\partial y}+\frac{\partial f}{\partial z}\]but this is not the gradient keep in mind operators have to be operating on something to be set equal to anything d/dx=k means nothing

OpenStudy (turingtest):

also remember that\[\nabla f\neq\nabla \cdot f\neq\nabla\times f\]and all that the middle is a scalar, the others are vectors

OpenStudy (anonymous):

What, what's the difference between ∇ƒ and ∇⋅ƒ

OpenStudy (anonymous):

And how can you take the cross product of the gradient operator and a function?

OpenStudy (turingtest):

a lot of difference\[\nabla f=<f_x,f_y,f_z>\]which is the gradiant, the vector perpendicular to the level curve of the function, as opposed to\[\nabla \cdot f=f_x+f_y+f_y\]which is a scalar you are not actually taking a dot product. that is impossible for the reason you stated. instead the notation of divergence is made similar to that of the dot product because of the obvious mathematical similarity it is a bit misleading though

OpenStudy (anonymous):

Ah, okay! Thanks!

OpenStudy (turingtest):

If we again think of F as the velocity field of a flowing fluid then div(F) represents the net rate of change of the mass of the fluid flowing from the point (x,y,z) per unit volume.

OpenStudy (turingtest):

welcome!